### v2

# SynFER: Towards Boosting Facial Expression Recognition with Synthetic Data

Abstract
Facial expression datasets remain limited in scale due to privacy concerns, the subjectivity of annotations, and the labor-intensive nature of data collection. This limitation poses a significant challenge for developing modern deep learning-based facial expression analysis models, particularly foundation models, that rely on large-scale data for optimal performance. To tackle the overarching and complex challenge, we introduce SynFER (Synthesis of Facial Expressions with Refined Control), a novel framework for synthesizing facial expression image data based on high-level textual descriptions as well as more fine-grained and precise control through facial action units. To ensure the quality and reliability of the synthetic data, we propose a semantic guidance technique to steer the generation process and a pseudo-label generator to help rectify the facial expression labels for the synthetic images. To demonstrate the generation fidelity and the effectiveness of the synthetic data from SynFER, we conduct extensive experiments on representation learning using both synthetic data and real-world data. Experiment results validate the efficacy of the proposed approach and the synthetic data. Notably, our approach achieves a 67.23% classification accuracy on AffectNet when training solely with synthetic data equivalent to the AffectNet training set size, which increases to 69.84% when scaling up to five times the original size. Our code will be made publicly available.
1 Introduction
Facial Expression Recognition (FER) is at the forefront of advancing AI’s ability to interpret human emotions, opening new frontiers for various human-centered applications. From automatic emotion detection to early interventions in mental health Ringeval et al. (2019), accurate pain assessment Huang et al. (2024), and enhancing human-computer interaction Abdat et al. (2011), the potential impact of FER systems is profound Moin et al. (2023); Sajjad et al. (2023); Zhu & Luo (2023). In recent years, learning-based FER models have gained significant traction due to their promising performances Li & Deng (2020); Zhang et al. (2021); Farzaneh & Qi (2021). However, despite recent advancements in network architectures and learning methodologies, the progress of existing FER models has been hindered by the inadequate scale and quality of available training data, underscoring the need to expand datasets with high-quality data to push the boundaries of FER capabilities.
Existing FER datasets, such as CK+ (953 sequences) Lucey et al. (2010), FER-2013 (30,000 48$\times$48 images) Barsoum et al. (2016), RAF-DB (29,672 images) Li et al. (2017), AFEW (113,355 images) Dhall et al. (2017), and SFEW (1,766 images) Dhall et al. (2011), are small compared to popular image datasets for general image processing (e.g., ImageNet Deng et al. (2009) with 1.4 million images and Laion Schuhmann et al. (2022) with billion-level data). While AffectNet Mollahosseini et al. (2017) compiles a large number of facial images from the web, it still suffers from vital drawbacks. A considerable portion of AffectNet’s images are low-quality, and its annotations often contain incorrect labels, which impairs the training process of FER models Le et al. (2023); Yan et al. (2022). Consequently, the absence of high-quality and large-scale FER datasets has delayed the development of FER foundation models. However, collecting a large-scale FER dataset with high-quality facial images and meticulous annotations is almost an unrealistic endeavor due to substantial financial and time costs, ethical concerns around facial data collection, and limited resources for large-scale acquisition. Additionally, the subjective interpretation of facial expressions results in inconsistent labeling by annotators, which exacerbates variability and hinders the creation of reliable datasets.
To address the challenges in developing FER models, we turn to synthesizing high-quality facial expression images paired with reliable labels. This approach draws inspiration from successful strategies employed to expand annotated datasets for other computer vision tasks, such as semantic segmentation Baranchuk et al. (2021); Chen et al. (2019); Li et al. (2022a) and depth estimation Atapour-Abarghouei & Breckon (2018); Cheng et al. (2020); Guizilini et al. (2022). These advances leverage powerful generative models such as Stable Diffusion Rombach et al. (2022a) and DALL-E Betker et al. (2023), which capture intricate natural image patterns. By tapping into these models, researchers have generated realistic images with their corresponding annotations, thereby boosting model performance. However, applying diffusion models to synthesize facial expression images with reliable FER labels presents two major challenges. (1) the training sets used by these generative models often lack diverse facial expression data, limiting their ability to produce images that capture subtle and nuanced emotional semantics; and (2) prior approaches to generate annotations for synthetic images focused on tangible attributes such as pixel-wise layouts, or depth maps. In contrast, facial expressions convey abstract and subjective emotions, making the generation of precise and reliable expression labels much more complex. To the best of our knowledge, none of the existing methods can simultaneously conduct fine-grained control for facial expression generation and generate robust categorical facial expression labels for face images.
Figure 1: (a) Examples of synthetic facial expression data generated by our SynFER model, (b) Comparison of training paradigms: training with real-world data versus training with synthetic facial expression data and (c) Performance boost from SynFER generated Data in supervised, self-supervised, zero-shot, and few-shot (5-shot) learning tasks.
In this paper, we present SynFER, the first framework capable of synthesizing unlimited, diverse and realistic facial expression images paired with reliable expression labels, to drive advancements in FER models. To address the shortcomings of existing FER datasets, which often lack expression-related text paired with facial images, we introduce FEText, a unique hybrid dataset created by curating and filtering data from existing FER and high-quality face datasets. This vision-language dataset serves as the foundation for training our generative model to synthesize facial expression data. To ensure fine-grained control and faithful generation of facial expression images, we inject facial action unit (FAU) information and semantic guidance from external pre-trained FER models. Building upon this, we propose FERAnno, the first diffusion-based label calibrator for FER, which automatically generates reliable annotations for the synthesized images. Together, these innovations position SynFER as a powerful tool for producing large-scale, high-quality facial expression data, offering a significant resource for the development of FER models.
We investigate the effectiveness of the synthetic data across various learning paradigms, demonstrating consistent and modest improvement in model performance. As shown in Fig. 1(c), training with the synthetic data yields significant performance boosts across various learning paradigms. Notably, pre-training on the synthetic data (Fig. 1(b)) with MoCo v3 Chen et al. (2021) yields a significant performance boost of +2.90% on AffectNet, surpassing real-world data pre-training. In supervised learning, SynFER improves accuracy by +1.55% for the state-of-the-art FER model, POSTER++ Mao et al. (2024), on AffectNet. We further explore performance scaling of the synthetic data, revealing further gains as dataset size increases.
Our key contributions are as follows: (1) we introduce FEText, the first dataset of facial expression-related image-text pairs, providing a crucial resource for advancing FER tasks. (2) we propose SynFER, the first diffusion-based data synthesis pipeline for FER, integrating FAUs information and semantic guidance to achieve fine-grained control and faithful expression generation. (3) we develop FERAnno, a novel diffusion-based label calibrator, designed to automatically refine and enhance the annotations of synthesized facial expression images. (4) extensive experiments across various datasets and learning paradigms demonstrate the effectiveness of the proposed SynFER, validating the quality and scalability of its synthesized FER data.
2 Related Work
Facial Expression Recognition (FER): Recent success in deep learning (DL) has largely boosted the performance of the FER task, despite the substantial data requirements for training DL models. To address the limited training data in FER, previous methods mainly focus on developing different learning paradigms, including semi-supervised learning Li et al. (2022b); Yu et al. (2023); Cho et al. (2024), transfer learning Li et al. (2022c); Ruan et al. (2022) and multi-task learning Liu et al. (2023b); Li et al. (2023a). For example, Ada-CM Li et al. (2022b) learns a confidence margin to make full use of the unlabeled facial expression data in a semi-supervised manner. Despite achieving performance gains for FER, these methods remain constrained by limited data. Recently, researchers have explored an alternative data-driven perspective of introducing large-scale face datasets from other facial analysis tasks (e.g., face recognition Zeng et al. (2022)). Meta-Face2Exp Zeng et al. (2022) utilizes large-scale face recognition data to enhance FER by matching the feature distribution between face recognition and FER. However, face data drawn from these datasets lack diverse facial expressions, and thereby couldn’t fully unlock the potential of large-scale data in FER.
Synthetic Data: Recently, growing attention has been paid to the advanced generative models (e.g., Generative Adversarial Networks (GANs) Goodfellow et al. (2020) and Diffusion Models Rombach et al. (2022b)), which are typically flexible to synthesize training images for a wider range of downstream tasks, including classification Frid-Adar et al. (2018); Azizi et al. (2023), face recognition Kim et al. (2023); Boutros et al. (2023), semantic segmentation Nguyen et al. (2023); Wu et al. (2024; 2023a) and human pose estimation Feng et al. (2023); Zhou et al. (2023). In particular, some studies pioneer to investigate the capabilities of powerful pre-trained diffusion generative models on natural images Nguyen et al. (2023); Wu et al. (2024); Li et al. (2023b). For example, DatasetDM Wu et al. (2024) further introduces a generalized perception decoder to parse the rich latent space of the pre-trained diffusion model for various downstream tasks. However, there remains a significant gap in research on using diffusion models to generate facial expression data, including both images and corresponding labels. In this paper, we address this gap by exploring diffusion-based synthetic data for the first time in FER. Specifically, we propose to apply a fine-tuned diffusion model to facial expression synthesis and introduce the first diffusion-based pseudo-label generator for FER.
3 Preliminaries
Diffusion models include a forward process that adds Gaussian noise $\epsilon$ to convert a clean sample $x_{0}$ to noise sample $x_{T}$, and a backward process that iteratively performs denoising from $x_{T}$ to $x_{0}$, where $T$ represents the total number of timesteps. The forward process of injecting noise can be formulated as:
$x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon$ (1)
$x_{t}$ is the noise feature at timestep $t$ and $\alpha_{t}$ is a predetermined hyperparameter for sampling $x_{t}$ with a given noise scheduler Song et al. (2020). In the backward process of denoising, given input noise $x_{t}$ sampled from a Gaussian distribution, a learnable network $\epsilon_{\theta}$ estimates the noise at each timestep $t$ with condition $c$. $x_{t-1}$, the feature at the previous timestep, is then derived as:
$x_{t-1}=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}x_{t}+\sqrt{\alpha_{t-1}}(\sqrt{\frac{1}{\alpha_{t-1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1})\epsilon_{\theta}(x_{t},t,c)$ (2)
During training, the noise estimation network $\epsilon_{\theta}$ is guided to conduct denoising with condition $c$ by the learning objective:
$\min_{\theta}\mathbb{E}_{x_{0},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),c,t}\|\epsilon-\epsilon_{\theta}(x_{t},c,t)\|_{2}^{2},$ (3)
With its powerful capability to model complex data distributions, the diffusion model serves as the foundation for generating high-quality FER data. Our SynFER framework is the pioneering work that explores the use of diffusion models to synthesize affective modalities.
4 Methodology
We begin by introducing i) the overall synthetic pipeline for generating facial expression image-label pairs. Next, we detail ii) our approach for producing high-fidelity facial expression images, which are controlled through high-level text descriptions (Sec.4.2.1), fine-grained facial action units corresponding to localized facial muscles (Sec.4.2.3), and a semantic guidance technique (Sec.4.3). Finally, we introduce iii) the FER annotation crafter (FERAnno), a crucial component that thoroughly understands the synthetic facial expression data and automatically generates accurate annotations accordingly(Sec.4.4). This pipeline ensures both precision and reliability in facial expression generation and labeling.
Figure 2: Overall pipeline of our FER data synthesis process.
4.1 Overall Pipeline for FER Data Synthesis
We introduce the overall pipeline for FER data synthesis (Fig. 2). The process starts with a coarse human portrait description assigned to a specific facial expression. ChatGPT enriches this description with details such as facial appearance, subtle facial muscle movements, and contextual cues. Simultaneously, facial action unit annotations are generated based on prior FAU-FE knowledge Ekman & Friesen (1978), aligning them with emotion categories to serve as explicit control signals for guiding the facial expression image synthesis. Once the facial expression label, facial action unit labels, and expanded textual prompt are prepared, these inputs condition our diffusion model to generate high-fidelity FER images, guided by semantic guidance to ensure accurate FER semantic. During the denoising process, FERAnno automatically produces pseudo labels for the generated images. To further improve labeling accuracy, we ensemble our FERAnno with existing FER models, which collaborate to vote on the accuracy of the predefined FER labels. In cases where discrepancies arise, the predefined label is refined by averaging the predictions from the ensemble experts. This mechanism effectively reduces the risk of inconsistent or uncertain annotations, ensuring that the final synthesis data is precise and dependable for downstream applications.
4.2 Diffusion Model Training for FER Data
4.2.1 FEText Data Construction
To address the lack of facial expression image-text pairs for diffusion model training, we introduce FEText (Fig. 3), the first hybrid image-text dataset for FER. It combines face images from FFHQ Karras et al. (2019), CelebA-HQ Karras et al. (2017), AffectNet Mollahosseini et al. (2017) and SFEW Dhall et al. (2011), each paired with captions generated by a multi-modal large language model (MLLM). FEText includes 400,000 curated pairs tailored for facial expression tasks.
Resolution Alignment. Due to variations in image resolution across different datasets, we first utilize a super-resolution model Lin et al. (2023) to standardize the resolutions of images from AffectNet and SFEW. Specifically, we incorporate high-resolution images from FFHQ and CelebA-HQ datasets to preserve the model’s capacity for high-fidelity image generation. This dual approach allows the model to not only maintain the fidelity of the generated images but also to learn and incorporate the facial expression semantics from AffectNet and SFEW.
Textual Caption Annotation. To generate a textural caption for each face image, we employ the open-source multi-modal language model ShareGPT-4V Chen et al. (2023b), by guiding it with carefully crafted instructions. To ensure that the generated captions are both context-aware and expressive, we clearly define the model’s role and provide examples of detailed facial expression descriptions within the prompts. This approach enables the model to generate precise, emotion-reflective captions for the input images.
4.2.2 Diffusion Model Fine-tuning
To facilitate our diffusion model to generate high-fidelity facial expressions, a straightforward approach is to fine-tune the model directly on the proposed FEText using the diffusion loss in Eq.  3. However, since FEText contains images processed through a super-resolution model, this direct fine-tuning strategy may lead to over-smoothing in the generated images. To address this, we introduce a two-stage fine-tuning paradigm. In the first stage, the diffusion model is trained on the entire FEText dataset to capture facial expression-related semantics. Then, the second stage mitigates over-smoothing by specifically fine-tuning our diffusion model on the CelebA-HQ and FFHQ subsets of FEText, which consist of native high-resolution images. This two-step approach ensures that our model learns expressive facial details while preserving image sharpness. The fine-tuned model then serves as the foundation for controllable facial expression generation, incorporating facial action unit injection (Sec. 4.2.3) and semantic guidance (Sec. 4.3).
Figure 3: Overview of our FEText data construction pipeline.
4.2.3 Explicit Control Signals via Facial Action Units
While fine-tuning the diffusion model using facial expression captions provides general language-based guidance for facial expression generation, it lacks the precision needed to capture fine-grained facial details, such as localized muscle movements. To overcome this limitation, we propose to incorporate more explicit control signals through Facial Action Units (FAUs), each of which represents a specific facial muscle movement. Inspired by IP-Adapter Ye et al. (2023), we apply a decoupled cross-attention module to integrate FAU embeddings with the diffusion model’s generation process. These embeddings are derived by mapping discrete FAU labels into high-dimensional representations using a Multi-Layer Perceptron, referred to as the AU adapter. FAU labels for each image in the FEText dataset are annotated using the widely adopted FAU detection model, OpenGraphAU Luo et al. (2022). With the diffusion model’s parameters frozen, we train the AU adapter to guide the model in recovering facial images based on the annotated FAU labels, using the objective in Eq. 3.
4.3 Semantic Guidance for Precise Expression Control
Due to the imbalanced distribution of FER labels in the training data and the potential ambiguity between certain facial expressions Zhang et al. (2024b), such as disgust, relying solely on textual and FAU conditions might not guarantee the faithful generation of these expressions. To address this issue, we propose incorporating semantic guidance on the textual embeddings $c^{\text{text}}$, during the later stages of the denoising process. We leverage external knowledge from open-source FER models to steer the generation process, ensuring a more accurate and faithful synthesis of hard-to-distinguish facial expressions.
Layout Initialization. During inference, we select a random face image $x^{s}$ from FEText and invert it to initialize the noise sample $x_{T}^{s}$(Eq. 1). Since early diffusion stages shape the global layout of the image Zhang et al. (2023); Pan et al. (2023); Mao et al. (2023), this strategy helps preserve the natural facial structure, ensuring the generated images are coherent, high-quality, and visually consistent with real-world expressions.
Semantic Guidance. In the early steps of the diffusion process, the generation process is conditioned on the original textual condition $c^{\text{text}}$. To further induce the generation of facial expression images corresponding to their FER labels $y$, we iteratively update the textual condition in the subsequent time steps. Specifically, a facial expression classifier $f(\cdot)$ is utilized for the injection of complex semantics.
To guide the generated images towards the specific class $y$, we propose to do so by updating the textual embeddings. Given an intermediate denoised sample $x_{t}$ at timestep $t$, following Eq. 15 in DDPM Ho et al. (2020), we first estimate the one-step prediction of the original image $\hat{x}_{0}$ as:
$\hat{x}_{0}=(x_{t}-\sqrt{1-\bar{\alpha}_{t}})\epsilon_{\theta}(x_{t},t,c^{\text{text}},c^{\text{au}})/\sqrt{\bar{\alpha}_{t}}$ (4)
We then calculate the classification loss with:
$\mathcal{L}_{g}=-y\log(h(f(\hat{x}_{0}))_{i})$ (5)
Given the guidance loss $\mathcal{L}_{g}$, the textual embedding is updated with the corresponding gradient:
$c^{\text{text}}_{t-1}=c^{\text{text}}_{t}+\lambda_{g}\frac{\nabla_{c^{\text{text}}_{t}}\mathcal{L}_{g}}{||\nabla_{c^{\text{text}}_{t}}\mathcal{L}_{g}||_{2}}$ (6)
where $\lambda_{g}$ and $c_{t-1}^{\text{text}}$ denote the step size and the updated textual embedding at timestep $t-1$, respectively. In the latter steps of the diffusion process, the noise estimator network $\epsilon_{\theta}$ is conditioned on the updated textual embeddings rather than the original one.
4.4 Diffusion-based Label Calibrator (FERAnno)
Figure 4: Overview of our FERAnno pseudo-label generator.
To ensure semantic alignment between each synthesized face image and its assigned facial expression label, we introduce FERAnno, a label calibration framework designed to validate the consistency of the generated data. By analyzing the facial patterns of each synthesized image, FERAnno categorizes them and compares the post-categorized labels with their pre-assigned facial expression labels. This verification process helps identify and filter out samples with mismatched labels, preventing them from negatively impacting downstream FER model training. Specifically, FERAnno is a diffusion-based label calibrator equipped with a deep understanding of facial semantics. It leverages the multi-scale intermediate features and cross-attention maps inherent in the diffusion model to predict accurate FER labels, as depicted in Fig. 4. This ensures only high-quality, correctly labeled samples are included in the training pipeline, leading to more reliable model performance.
Image Inversion. To extract facial features and cross-attention maps with the diffusion model $\epsilon_{\theta}$, we first inverse the generated image $x_{0}$ back to the noise sample $x_{t}$ at a denoising timestep $t$, following a predefined scheduler, as described in Eq. 1. To preserve facial details, we set $t=1$ during the inversion process, ensuring that the facial features remain as close as possible to the original generated image $x_{0}$. This partially denoised sample is then passed through the trained denoising network, allowing us to extract rich facial features and cross-attention maps from intermediate layers, which are critical for capturing detailed facial patterns.
Feature Extraction. Given the inverted noise sample $x_{1}$ and the corresponding textual condition $c^{\text{text}}$ and AU condition $c^{\text{au}}$, we can extract the multi-scale feature representations and textual cross-attention maps from the U-Net $\epsilon_{\theta}$ as $\left\{\mathcal{F},\mathcal{A}\right\}=\epsilon_{\theta}(x_{1},t_{1},c^{\text{text}},c^{\text{au}})$, where $\mathcal{F}$ and $\mathcal{A}$ denote the multi-scale feature representations and the cross-attention maps, respectively. $\mathcal{F}$ contains multi-scale feature maps from different layers of the U-Net $\epsilon_{\theta}$ with four different resolutions. $\mathcal{A}$ contains the cross-attention maps drawn from the 16 cross-attention blocks in $\epsilon_{\theta}$. Both the feature representation $\mathcal{F}$ and the cross-attention maps $\mathcal{A}$ are regrouped according to their resolutions.
Multi-scale Features and Attention Maps Fusion. Given that the multi-scale feature maps $\mathcal{F}$ capture global information essential for image generation, and the cross-attention maps provide class-discriminative information as well as relationships between object locations Tang et al. (2022); Caron et al. (2021), FERAnno fuses both features and attention maps within a dual-branch encoder architecture for pseudo-label annotation. An overview of this architecture is shown in Fig. 4.
We first compute the mean of the regrouped attention maps, denoted as $\mathcal{A}_{\text{reg}}$, yielding a set of averaged attention maps $\bar{\mathcal{A}}$. Both the feature maps $\mathcal{F}$ and the averaged attention maps $\bar{\mathcal{A}}$ are then passed through a residual convolution block to prepare them for further processing. To effectively integrate information at different scales, we introduce a bi-directional cross-attention block to fuse the features and attention maps. 1$\times$1 convolutions are employed at various stages to adapt the fusion across multiple resolution layers. Finally, the fused feature maps and attention maps are concatenated and passed through a linear layer, which outputs a probability vector for predicting facial expression classes.
5 Experiments
We conduct extensive experiments to evaluate both the generation quality of our synthetic data (Sec. 5.1) and its effectiveness in FER tasks (Sec. 5.2). For more details on experimental setup, implementation details are provided in the appendix.
5.1 Generation Quality
Method Objective Metrics User study (Ours vs. )(%)
FID ($\downarrow$) HPSv2($\uparrow$) FS($\uparrow$) MPS ($\uparrow$) FER Acc.($\uparrow$) FAU Acc.($\uparrow$) EA ($\uparrow$) FF ($\uparrow$)
Stable Diffusion 88.40 0.263 2.01 2.00 20.06 87.72 2.86 1.79
PixelArt 145.23 0.271 3.79 5.26 15.52 84.57 24.26 10.00
PlayGround 81.76 0.265 2.86 3.73 21.56 87.28 7.50 5.00
FineFace 74.61 0.268 3.29 1.48 38.05 89.68 5.73 6.41
SynFER 16.32 0.280 4.26 - 55.14 93.31 59.64 76.79
Table 1: Ours vs.’ shows the proportion of users who prefer our method over the alternative. An MPS above 1.00 and results above 50% in the user study indicate our method outplays the counterpart. FS, FER Acc., FAU Acc., EA and FF denote FaceScore Liao et al. (2024), FER accuracy, facial action unit accuracy, expression alignment and face fidelity, respectively.
We present both objective metrics and subjective user studies, comparing our method to existing state-of-the-art (SOTA) diffusion models Rombach et al. (2022a); Chen et al. (2023a); Li et al. (2024) and the latest facial expression generation technique, FineFace Varanka et al. (2024).
We compute FID between the synthesis images and the test set of the AffectNet Mollahosseini et al. (2017). HPSv2 Wu et al. (2023b) and MPS Zhang et al. (2024a) evaluate the human preferences of the overall synthesis images, while FaceScore Liao et al. (2024) measures the quality of the generated faces. Tab. 1 shows that our method outperforms popular diffusion models and the SOTA facial expression generation method FineFace Varanka et al. (2024), across all metrics of image quality, human preference and facial expression accuracy. Notably, the advantages of SynFER in both FE Acc. and AU Acc. indicate its outstanding controllability in facial expression generation.
5.2 Effectiveness of Synthetic Dataset
Method Pre-train Data RAF-DB AffectNet SFEW
Dataset Scale
MCF Laion-Face 20M 65.22 - 32.61
FRA VGGFace2 3.3M 73.89 57.38 -
PCL VoxCeleb 1.8M 74.47 - 39.68
SimCLR AffectNet 0.2M 78.65 48.36 46.79
SimCLR Ours 1.0M 80.24 (+1.59) 52.05 (+3.69) 47.62 (+0.83)
SimCLR AffectNet+Ours 1.2M 81.52 (+2.87) 54.37 (+6.01) 48.52 (+1.73)
BYOL AffectNet 0.2M 78.24 50.04 48.70
BYOL Ours 1.0M 80.96 (+2.72) 53.13 (+3.09) 51.35 (+2.65)
BYOL AffectNet+Ours 1.2M 81.25 (+3.01) 54.95 (+4.91) 51.70 (+3.00)
MoCo v3 AffectNet 0.2M 79.05 51.03 49.34
MoCo v3 Ours 1.0M 81.17 (+2.12) 55.56(+4.53) 50.78 (+1.44)
MoCo v3 AffectNet+Ours 1.2M 81.68 (+2.63) 57.84 (+6.81) 51.26 (+1.92)
Table 2: Linear probe performance comparisons of SSL models on three FER datasets.
Method RAF-DB AffectNet
ResNet-18 87.48 50.32
ResNet-18 + Ours 87.97 51.65
Ada-DF 90.94 65.34
Ada-DF + Ours 91.21 66.82
POSTER++ 91.59 67.49
POSTER++ + Ours 91.95 69.04
APViT 91.78 66.94
APViT + Ours 92.05 67.26
FERAnno 92.56 70.38
Table 3: Comparison of supervised learning models (with and without our synthetic data) and the label calibrator FERAnno.
Self-supervised Representation Learning. We trained self-supervised learning (SSL) models, including BYOL Grill et al. (2020), MoCo v3 Chen et al. (2021), and SimCLR Chen et al. (2020), using real-world data, our synthetic data, and a combination of both. The linear probe performance of these models was evaluated on three widely used facial expression recognition (FER) datasets: RAF-DB Li et al. (2017), AffectNet Mollahosseini et al. (2017), and SFEW Dhall et al. (2011), with results reported in Tab. 3. All SSL models were trained with a ResNet-50 architecture He et al. (2016). Notably, state-of-the-art methods in self-supervised facial representation learning, such as MCF Wang et al. (2023), FRA Gao & Patras (2024), and PCL Liu et al. (2023c), were pre-trained on much larger face datasets like LAION-Face Zheng et al. (2022), VGGFace2 Cao et al. (2018), and VoxCeleb Nagrani et al. (2020). However, these models underperformed on FER tasks compared to ours, highlighting that existing large-scale face datasets may lack the high-quality and diverse facial expression patterns required for accurate FER. Results demonstrate that combining real-world and synthetic data consistently boosts SSL baselines. Remarkably, even when MoCo v3 was trained solely on our synthetic data, it achieved a 2.12% improvement on RAF-DB, underscoring the effectiveness of our approach in capturing critical facial expression details that are essential for FER.
Supervised Representation Learning.
Method CFEE_C EmotionNet_C RAF_C
1-shot 5-shot 1-shot 5-shot 1-shot 5-shot
InfoPatch 54.19 67.29 48.14 59.84 41.02 57.98
InfoPatch* 55.21 68.73 48.52 61.16 41.88 59.54
LR+DC 53.20 64.18 52.09 60.12 42.90 56.74
LR+DC* 54.65 65.28 51.96 60.14 43.87 57.90
STARTUP 54.89 67.79 52.61 61.95 43.97 59.14
STARTUP* 56.25 69.93 52.87 62.12 45.18 61.23
CDNet 56.99 68.98 55.16 63.03 46.07 63.03
CDNet* 57.74 70.64 56.79 65.63 46.97 64.34
Table 4: Performance comparisons with SOTA few-shot learning methods on 5-way few-shot FER tasks with a 95% confidence interval. (*) indicates training with both real-world data and our synthesis data.
Method RAF-DB AffectNet FERPlus
UAR WAR UAR WAR UAR WAR
FaRL 24.98 38.53 26.95 26.95 24.27 35.26
FLAVA 14.35 38.69 14.26 14.26 12.50 28.43
CLIP ViT-B/16 38.66 36.16 34.35 34.35 34.33 45.14
Exp-CLIP ViT-B/16 48.96 54.50 39.98 39.98 40.81 53.02
Exp-CLIP* ViT-B/16 51.23 56.32 41.68 41.68 42.33 56.02
CLIP ViT-L/14 47.22 41.13 34.46 34.47 33.82 46.67
Exp-CLIP ViT-L/14 58.70 65.37 44.27 44.27 48.28 55.42
Exp-CLIP* ViT-L/14 60.41 68.34 46.85 46.85 50.26 57.98
Table 5: Zero-shot performance comparison of CLIP for FER, reporting both Weighted Average Recall (WAR) and Unweighted Average Recall (UAR) as in previous works Zhao et al. (2024). (*) indicates models trained with both real-world and synthetic data.
We validate the effectiveness of SynFER for supervised representation learning by evaluating its performance on RAF-DB and AffectNet (Tab. 3). We compare with SOTA FER models, including Ada-DF Liu et al. (2023a), POSTER++ Mao et al. (2024), and APViT Xue et al. (2022). The results demonstrate that incorporating synthetic data consistently enhances both baseline models and the latest SOTAs in supervised facial expression recognition. Notably, APViT benefits from the synthetic data with improvements of 0.27% on RAF-DB and 0.32% on AffectNet. While the improvements in supervised learning are more modest compared to self-supervised learning, they remain consistent. This is likely due to the stricter distribution alignment required in supervised learning between synthetic training data and real-world test data. In the following section on scaling behavior analysis, we provide further insights, showcasing the use of the distribution alignment technique, Real-Fake Yuan et al. (2023), to alleviate this problem.
Few-shot Learning. Addressing the challenge of limited labeled FER data across different scenarios, we explore the potential of synthetic data to enhance few-shot learning, as presented in Tab. 5. Following the protocol established by CDNet Zou et al. (2022), we train models on five basic expression datasets and evaluate them on three compound expression datasets: CFEE_C Du et al. (2014), EmotionNet_C Fabian Benitez-Quiroz et al. (2016), and RAF_C Li et al. (2017). To benchmark our approach, we compare it against SOTA few-shot learning methods, including InfoPatch Liu et al. (2021), LR+DC Yang et al. (2021), and STARTUP Phoo & Hariharan (2021). The results clearly demonstrate that integrating synthetic data consistently enhances few-shot FER performance across key metrics. This highlights the ability of synthetic data, with its broader range of FER patterns, to bridge the gap in data-limited scenarios, allowing models to better generalize to complex, real-world expressions in few-shot tasks.
Multi-modal Fine-tuning. The synthesis data encompasses multiple modalities, including generated images, textual prompts, and FER labels. To assess its impact on multi-modal fine-tuning for FER, we focus on fine-tuning the vision-language foundation model CLIP Wang et al. (2022), as its performance on face-related tasks is widely regarded as sub-optimal Guo et al. (2023); Chen et al. (2024). Building on Exp-CLIP Zhao et al. (2024), we fine-tune the models on CAER-S Lee et al. (2019) and evaluate their zero-shot performances on the datasets outlined in Tab. 5. Our results show that the inclusion of detailed textual prompts and a larger training image set significantly enhances the generalization ability of Exp-CLIP in understanding facial expressions, achieving significant improvements such as +2.58% UAR on the AffectNet dataset.
5.3 Ablation Study
Effectiveness of FAU Control. We validate the effectiveness of SynFER by examining how Facial Action Units (FAUs) enhance the generation process and refine facial details. As illustrated in Fig. 5, samples generated with FAU control (third column) exhibit facial expressions that more accurately match their assigned labels compared to those generated with only text guidance (second column). For example, the ’fear’ expression, driven by FAUs like Inner Brow Raiser and Lip Stretcher, becomes more distinct (third column, second row), making it easier to differentiate from other emotions such as ’surprise.’ Similarly, ’disgust’ is more pronounced with FAUs like Lid Tightener. Without FAU control, facial expressions (second column) tend to blur, as different categories show overlapping features. Quantitative results in Tab. 6 highlight the impact of FAU control: FER accuracy increases from 34.62% to 48.74%, and FAU detection accuracy rises from 88.91% to 92.37%. This also translates into improved downstream performance on RAF-DB and AffectNet.
Figure 5: Generated samples. The first and second rows are fear and disgust, respectively.
Method HPSv2 FE Acc. AU Acc. RAF-DB AffectNet
Real-world Data - - - 91.59 67.49
SD 0.263 20.06 87.72 89.42 65.36
w/ FEText 0.267 34.62 88.91 90.54 66.62
w/ FEText+FAUs 0.275 48.74 92.37 91.68 67.68
w/ FEText+FAUs+SG 0.280 55.14 93.31 91.95 68.13
Table 6: Ablation study on the influence of AU injection and semantic guidance (SG) on both the generation quality and supervised representation learning. SD denotes Stable Diffusion , which is used as a baseline.
Effectiveness of Semantic Guidance. We further explore the impact of semantic guidance (SG) on both generation quality and supervised representation learning, as shown in Fig. 5 and Tab. 6. By updating text embeddings to better align with the target facial expression category, SG improves the accuracy of the generated expressions by 6.4%, compared to static text and FAUs. The samples in the last column of Fig. 5 show more exaggerated facial expressions than those in the third column, with SG enhancing the intensity.
Reliability of FERAnno. We assess the reliability of FERAnno as a label calibrator by evaluating its performance on two FER datasets and visualizing its attention maps in Tab. 3 and Fig. 6. FERAnno significantly outperforms previous SOTAs, achieving a +0.51% improvement on RAF-DB and a +1.34% improvement on AffectNet over the second-best models. The attention maps in Fig. 6 further demonstrate FERAnno’s ability to accurately locate facial expression-related facial features, such as jaw-dropping and furrowed eyebrows, highlighting the diffusion model’s great semantic understanding and fine-grained facial expression recognition.
Figure 6: Synthesis images and attention maps in the fine-tuned diffusion model.
Synthetic Data Scaling Analysis. Following Tian et al. (2024); Fan et al. (2024), we investigate the scaling behavior of synthetic data in both self-supervised and supervised learning paradigms. To highlight the potential of synthetic FER data, we train models exclusively on synthetic images, without combining real-world data. The results in Fig. 7 (a)-(b) show a stronger scaling effect in self-supervised learning compared to supervised learning, where performance improves significantly with more data. This difference is likely due to the need for better distribution alignment in supervised learning Yuan et al. (2023). While SynFER focuses on addressing FER data scarcity, aligning the synthetic data distribution with real-world data is crucial for supervised tasks. To further explore this, we apply the Real-Fake technique Yuan et al. (2023) for real and synthetic data distribution alignment, and present the results in Fig. 7 (c). Compared to standard supervised learning, Real-Fake demonstrates a clear performance boost.
Figure 7: Scaling up the synthetic FER dataset. MoCo v3 (ResNet-50) Chen et al. (2021) is used for SSL pre-training, and linear probe performance is evaluated on AffectNet and RAF-DB. The SOTA FER model, POSTER++ Mao et al. (2024), is trained using supervised learning (with and without the Real-Fake technique Yuan et al. (2023)) on our synthetic dataset and evaluated on the same two target FER datasets. ★ is model’s performance trained on corresponding real data.
6 Conclusion
In this paper, we propose a synthesis data framework SynFER for facial expression recognition to address the data shortage in the field. By consolidating existing FER data and annotating with multi-modal large language models, we introduce the first facial expression-related image-text pair hybrid dataset FEText. We further propose to inject facial action unit information and external knowledge from existing FER models to ensure both fine-grained control and faithful generation of the facial expression images. To incorporate the generated images into training, we propose a diffusion-based label calibrator to help rectify the robust facial expression annotations for the synthesized images. After constructing the data synthesis methodology, we investigate the effectiveness of the synthesis data across different learning paradigms, demonstrating consistent and superior performances. We further study the scaling behavior of the synthesis data for FER.
Appendix A Appendix
A.1 Hyper-parameter Studies
Step Size $\lambda$ in Semantic Guidance. For hyper-parameter analysis, we consider five configurations of the step size $\lambda$ in semantic guidance. Due to computational resource constraints, we provide results of self-supervised learning with MoCo v3 Chen et al. (2021) on 0.2M synthetic data for pre-training and report the linear probe performances on RAF-DB Li et al. (2017). Experiment results are shown in Fig. 8. It can be seen that when $\lambda$ is relatively small, its influence on the performance is relatively small. However, as $\lambda$ continues to increase, the downstream performance is severely degraded. This is because an excessive $\lambda$ would lead to severely disrupted images, as shown in Fig. 9.
Figure 8: Hyper-parameter study on the $\lambda$ in semantic guidance. We report the linear probe performances of MoCo v3 Chen et al. (2021) pre-trained with 0.2M synthetic data on RAF-DB Li et al. (2017).
Figure 9: Visualizations of images generated with different values of $\lambda$.
A.2 Experiment Setting and Implementation Details
Self-Supervised Learning. We use the widely adopted self-supervised learning library solo-learn Da Costa et al. (2022) for experiments and follow the default settings in solo-learn for various methods. Detailed settings are shown in the tables below:
Config Pre-Training Linear Probe
SimCLR BYOL MoCo v3 SimCLR BYOL MoCo v3
batch size 64 64 64 32 32 32
optimizer Lars Lars Lars SGD SGD SGD
base learning rate 0.3 0.1 0.3 1e-3 1e-3 1e-3
weight decay 1e-4 1e-6 1e-6 1e-4 1e-4 1e-4
learning rate schedule warmup cosine warmup cosine warmup cosine step (60,80) steps (60,80) steps (60,80)
epochs 200 200 200 100 100 100
augmentation RRC RRC RRC RRC+RHF RRC+RHF RRC+RHF
Table 7: Implementation details on self-supervised pre-training. RRC and RHF denote random resize crop and random horizontal flip, respectively.
Others. As all the methods for comparisons in supervised learning (Tab. 3), few-shot learning (Tab. 5) and multi-modal fine-tuning (Tab. 5) are open-source, we thus only need to rewrite the corresponding code for dataset reading to incorporate the synthetic data. We follow the default setting in each open-source code of the compared methods.
Facial Action Unit Setting. We use pre-defined facial action unit (FAU) labels to generate images corresponding to specific facial expressions as shown below:
Facial Expression FAU
Happy AU6 + AU12
Sad AU1 + AU4 + AU15
Surprise AU1 + AU2 + AU5 + AU26
Fear AU1 + AU2 + AU4 + AU5 + AU7 + AU20 + AU26
Angry AU4 + AU5 + AU7 + AU23
Disgust AU9 + AU15 + AU16
Table 8: FAU annotations to generate specific classes of facial expression images.
Step of Performing Semantic Guidance. During the synthesis process, the total denoising steps of the diffusion model are set as 50. Semantic guidance requires backward gradient computation, which would cost a large amount of GPU hours. Thereby, we only perform semantic guidance in the latter steps, which is set as the last 5 steps of the denoising. Another reason to perform semantic guidance at the latter steps is that estimated results at early steps tend to be blurry and degraded facial images, performing semantic guidance on such images might to incorrect results.
Figure 10: Synthetic images comparison between the over-smoothing images and the natural images.
A.3 Over-smoothing of Super Resolution Training Data
As shown in Fig. 10, we provide comparisons between the over-smoothing synthetic images and the more natural ones. Due to the large amount of super-resolution data in FEText, it can be seen that solely performing fine-tuning on the entire FEText significantly degrades the realism of the images, while the proposed two-stage fine-tuning strategies in Sec. 4.2.2 could effectively prevent over-smoothing.
A.4 Limitations and future work
While the effectiveness of the proposed synthetic data framework has been demonstrated through extensive experiments, its current use is limited to augmenting the training set. A more efficient and optimized approach for leveraging synthetic data remains unexplored and warrants further investigation. Additionally, the generation process remains relatively slow, particularly when incorporating semantic guidance, which is crucial for ensuring accurate and faithful data generation. Moreover, this work focuses exclusively on facial expression recognition. However, it is important to note that the synthetic data framework has potential applications in other areas of facial affective computing, such as facial action unit detection and affective valence and arousal recognition. These avenues are left for future exploration.
A.5 FEText
More examples from FEText are shown in Fig. 11.
Figure 11: Examples from FEText.